# SOMA: a novel sampler for exchangeable variables

### v1

@@ 3 SOMA: A Single-Offer-Multiple-Attempts sampler
SOMA sampler with target distribution $\pi$.
     Input: proposal kernel $q(\cdot)$, number of SOMA iteration $T$
     Initialization: set initial value of $x=(x_{1},\ldots,x_{n})^{\top}$.
     For $i=1$ to $i=T$
@@
        Propose $y\sim q(\cdot)$
        Set weight $w_{0}=\frac{\pi(x)}{\prod_{j=1}^{n}q(x_{j})}$, $w_{i}=\frac{\pi([x_{-i},y])}{q(y)\prod_{j=1,j\neq i}^{n}q(x_{j})}$ and $W=\sum_{i=1}^{n}w_{i}$
        Select random index $I\in\{1,2,\ldots,n\}$ with probability $\mathbb{P}(I=i)=\frac{w_{i}}{W}$
      Stage 2:
        Replace $x_{I}$ with $y$ using probability $\alpha=\min\left\{1,\frac{W}{W+w_{0}-w_{I}}\right\}$
        Update state $x^{(t)}\leftarrow(x_{1},\ldots,x_{n})^{\top}$
     Output: SOMA samples $\{x^{(t)}\}_{t=1}^{T}$
While IMwG exploits the symmetry of permutation invariant distributions by using the same independent proposal for each component, SOMA has pushed this idea further. SOMA recognizes that an unfavorable proposal for $x_{j}$ might be a favorable proposal for some other $x_{i}$ and it seeks to find such an index $i$ by comparing all the weights given in (5).
The advantage of SOMA lies in its efficient use of the ‘single offer’ from the proposal distribution $q$. Stage 1 of SOMA chooses a favorable component $i$ among the multiple choices to swap out $x_{i}$ by $y$. Then Stage 2 compares the current state $x$ with the proposed state $x^{\prime}=[x_{-i},y]$. The acceptance probability (6) accounts for the fact that $x^{\prime}$ is selected from multiple attempts and is not a direct comparison like (4). We summarize SOMA in Algorithm 1 and show that it is reversible with respect to the target $\pi$ in Theorem 2. The proof is deferred to Section B.1.
@@
Assume $\pi$ is permutation invariant. Stage 1 of Algorithm 1 generates a proposal kernel, which we denote by $K^{\mathrm{SOMA}}(x^{\prime}\mid x)$. Let $d(x,x^{\prime})$ be the minimum number of components in which $x$ and $x^{\prime}$ differ, up to permutation of indices, i.e.,
$d(x,x^{\prime})=\min_{\sigma\in S_{n}}\|x_{\sigma}-x^{\prime}\|_{0},$
where we use $x_{\sigma}$ to represent the vector obtained from permitting indices from $x$. We must have $K^{\mathrm{SOMA}}(x^{\prime}\mid x)=0$ for all $d(x,x^{\prime})>1$. When $d(x,x^{\prime})=1$, suppose they only differ on the $i$-th entry, i.e., $x_{i}^{\prime}=y$ and $x_{j}^{\prime}=x_{j}$ for all $j\neq i$, then we must have $K^{\mathrm{SOMA}}(x^{\prime}\mid x)=q(y){w_{i}(y,x)}/{W(y,x)}.$ Associating this proposal kernel with acceptance probability (6) in Stage 2 leads to a Markov transition kernel $P_{\mathrm{SOMA}}$ that is reversible with respect to $\pi$.
@@
From Theorem 2, we see that SOMA transforms the independent proposal $q$ into a proposal kernel $K^{\mathrm{SOMA}}$ that is dependent on the current state. This distinguishes SOMA from IMH. The proposal kernel $K^{\mathrm{SOMA}}$ is also different in nature from common MCMC algorithms such as random walk Metropolis, Hamiltonian Monte Carlo, and Langevin Monte Carlo. SOMA does not require gradient information or a step-size.
@@
We should also distinguish the SOMA sampler from the Ran-IMwG and Sys-IMwG samplers. To begin with, SOMA is not a Gibbs sampler - it is a Metropolis-Hastings type algorithm, consisting of a proposal step and an accept/reject step. Instead, Stage 1 of SOMA constructs a Metropolis kernel $K^{\mathrm{SOMA}}$ by converting the IMH proposal into a dependent one. Second, although SOMA and Ran-IMwG randomly pick an index $i$ to perform a swap move, Ran-IMwG uses fixed weights independent of the current state, while SOMA uses weights (5) that adapt to the current state. We can see SOMA as a multiproposal MCMC (Pozza and Zanella,, 2025) method, and Ran-IMwG is the ‘single proposal’ counterpart that SOMA should compare with. More discussions are in Section A.
@@
Multiple-try Metropolis (MTM) (Liu et al.,, 2000) also entails multiple evaluations in one Metropolis iteration. In particular, the Multiple-try Metropolis Independent sampler (MTM-IS) (Yang and Liu,, 2023) bears a resemblance to SOMA. MTM-IS generates multiple samples $\{y_{1},\ldots,y_{n}\}$ from an independent proposal $q$, selects one index from them, and generates ‘balancing trials’ to satisfy reversibility. On the other hand, SOMA only requires one sample $y$ from $q$. The multiple attempts for SOMA are created by comparing $y$ with each component in the current state $x$. SOMA would require much less computation and memory cost than MTM-IS.
@@ 4.1 Acceptance rate
For accept-reject-based Markov chains, a small acceptance probability is a clear signal for slow convergence (Brown and Jones,, 2022). For component-wise Markov chains, slow convergence of the Metropolis kernels makes the overall Metropolis within Gibbs sampler mix slowly (Qin et al.,, 2023).
First, we analyze the coordinate-wise acceptance probability of the three algorithms. This is the probability of accepting a proposed move conditioning on the proposal point as well as the chosen coordinate for a swap move.
@@
The complete proof is in Section B.2. Theorem 4 indicates that, if all three methods choose to update the same component $i$, SOMA always has an acceptance probability at least as high as the other IMwG methods.
@@
  • Proof.
    Suppose $x$ and $x^{\prime}$ differs on component $i$ and with out loss of generality assume $i=1$. Then, starting from the right-hand side,
    $n\cdot P_{\mathrm{RAN}}(x^{\prime}\mid x)$ $=n\cdot\frac{1}{n}q(x^{\prime}_{1})\alpha_{1}^{\mathrm{IMwG}}(x^{\prime}_{1},x)=n\cdot\frac{1}{n}q(x^{\prime}_{1})\min\left(\frac{w_{1}}{w_{0}},1\right)$
    $=q(x^{\prime}_{1})\cdot w_{1}\cdot\min\left(\frac{1}{w_{1}},\frac{1}{w_{0}}\right)$
    $\geq q(x^{\prime}_{1})\cdot w_{1}\cdot\min\left(\frac{1}{w_{1}+\sum_{i=2}^{n}w_{i}},\frac{1}{w_{0}+\sum_{i=2}^{n}w_{i}}\right)$
    $=q(x^{\prime}_{1})\cdot\frac{w_{1}}{w_{1}+\sum_{i=2}^{n}w_{i}}\min\left(1,\frac{w_{1}+\sum_{i=2}^{n}w_{i}}{w_{0}+\sum_{i=2}^{n}w_{i}}\right)=P_{\mathrm{SOMA}}(x^{\prime}\mid x).$
@@
Next, we analyze the overall acceptance probabilities. Let $A(x)$ be the probability of accepting a move from $x$. For SOMA and Ran-IMwG, the coordinate $i$ is chosen randomly and we have
@@
To establish lower bounds for the overall acceptance probability $A(x)$ for general $n$, we make the following assumption about the proposal kernel $q$ and the target distribution $\pi$.
Assumption 2 (Bounded ratio of weights).
For all indices $1\leq i,j\leq n$, the ratio between the weights $w_{i}(y,x)$ and $w_{j}(y,x)$ is bounded, i.e., there exists a constant $M>1$ such that
$\frac{w_{i}(y,x)}{w_{j}(y,x)}\leq M\quad\forall\quad 1\leq i,j\leq n,x\in\mathbb{X}^{n},y\in\mathbb{X}.$ (16)
In other words, the proposal distribution $q$ and target distribution $\pi$ satisfies
$\frac{\pi(y,x_{-i})q(x_{i})}{\pi(y,x_{-j})q(x_{j})}\leq M\quad\forall\quad 1\leq i,j\leq n,x\in\mathbb{X}^{n},y\in\mathbb{X}.$ (17)
A sufficient condition for Assumption 2 to hold is for the weights to be bounded, i.e., there exist constants $M_{2}>M_{1}>0$, such that $M_{1}\leq w_{i}(y,x)\leq M_{2}$ for all $i,y,x$. Bounded weights have been a common assumption in Monte Carlo methods. For example, for single-component IMH updates (i.e., $n=1$), a uniformly bounded ratio $\pi(x)/q(x)\leq M$ for all $x\in\mathbb{X}$ is required to ensure geometric ergodicity (Mengersen and Tweedie,, 1996; Wang,, 2022; Yang and Liu,, 2023).
We give instances where Assumption 2 holds.
Definition 8 ($\epsilon$-differential privacy ($\epsilon$-DP) (Dwork,, 2006)).
We say a randomized algorithm with conditional density function $\eta(s_{\mathrm{dp}}\mid x)$ satisfies if $\epsilon-$DP if, for all $s_{\mathrm{dp}}$ and for all pairs of neighboring databases $(x,x^{\prime})$ (i.e., $d(x,x^{\prime})=1$), we have
$\frac{\eta(s_{\mathrm{dp}}\mid x)}{\eta(s_{\mathrm{dp}}\mid x^{\prime})}\leq\exp(\epsilon),$ (18)
where the parameter $\epsilon>0$ is called the privacy loss budget.
Let’s consider data imputation after observing summary statistics through an $\epsilon$-DP noise. Revisiting Example 2, the target distribution (2) has joint density $\pi(x\mid s_{\mathrm{dp}})\propto\eta(s_{\mathrm{dp}}\mid x)\prod_{i=1}^{n}f(x_{i})$.
Example 4 (Data imputation given privatized summary).
Let’s assume $\eta$ satisfied $\epsilon$-DP. Suppose we choose the proposal distribution $q$ to be $f$ in Algorithm 1. The weights are $w_{i}(y,x)\propto\eta(s_{\mathrm{dp}}\mid y,x_{-i})$ and $w_{0}(y,x)\propto\eta(s_{\mathrm{dp}}\mid x)$, and Assumption 2 holds with $M=\exp(\epsilon)$.
We now give lower bounds for the overall acceptance probability from the current state $x$.
@@
Let $A^{\mathrm{SOMA}}(x)$, $A^{\mathrm{RAN}}(x)$ and $A_{i}^{\mathrm{SYS}}(x)$ be defined as in (11), (12) and (13) respectively. These are the overall acceptance probabilities from a current state $x$. With Assumption 2, we have
$A^{\mathrm{SOMA}}(x)\geq\frac{n}{n+M-1},\ A^{\mathrm{RAN}}(x)\geq\frac{1}{M},\ \text{and}\ A_{i}^{\mathrm{SYS}}(x)\geq\frac{1}{M}\ \forall i.$ (19)
The lower bounds on acceptance probability place the three algorithms into a ‘no clear signal of slow convergence’ scenario. We also highlight that when $n$ is large, the acceptance probability of SOMA approaches 1; which we confirm with experiments in Section 5. As a result, SOMA becomes almost rejection-free in high dimensions.
@@ 4.2 Upper bound of the convergence rate
In this section, we use coupling-based techniques to analyze the convergence rate of the algorithms. Our proof leverages the coupling lemma. It entails constructing two chains on a common probability space such that they eventually meet and remain coupled.
Theorem 10 (Coupling inequality (Rosenthal,, 1995)).
$\|P^{t}(\cdot\mid x)-\pi(\cdot)\|_{\mathrm{TV}}\leq\mathbb{P}(\tau>t),\quad\text{where meeting time }\tau=\min\{t:\ \Phi^{(t)}=\tilde{\Phi}^{(t)}\}.$ (20)
To utilize (20), we must construct such a faithful coupling. We create a coupling of SOMA in the two stages: (1) use a common proposed state $y$ and select a common index $I$ using a maximal coupling of Multinomial random variables (Thorisson,, 2000; Jacob,, 2020); (2) accept and reject the proposed states with a common random number. We summarize the coupling kernel in Algorithm 2.
Algorithm 2.
One iteration of a coupling kernel for SOMA
  1. 0.
    Input: current states $x$ and $\tilde{x}$ from $\Phi$ and $\tilde{\Phi}$
  2. 1.
    Propose $y\sim q(y)$ as common single proposal
  3. 2.
    Calculate weights $w_{i}$ and $\tilde{w}_{i}$ according to (5) and (7)
  4. 3.
    Select components $I$ and $\tilde{I}$ via maximal coupling of $\mathrm{Multinomial}(w)$ and $\mathrm{Multinomial}(\tilde{w})$
  5. 4.
    Compute probabilities $\alpha_{I}$ and $\tilde{\alpha}_{\tilde{I}}$ according to (6) and use a common random number $U\sim\mathrm{Uniform}(0,1)$ to accept/reject each chain.
We include details of maximal coupling of multinomial distributions in the Supplementary Materials for completeness. We provide the following property of Algorithm 2.
Lemma 11.
For any current state $(x,\tilde{x})$, let $(x^{\prime},\tilde{x}^{\prime})$ be samples from the coupled SOMA kernel $P_{\mathrm{SOMA-coupled}}$ in Algorithm 2. Under Assumption 2, we must have
$P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})-1\mid x,\tilde{x}\right)\geq\frac{n}{n+M-1}\cdot\frac{d(x,\tilde{x})}{d(x,\tilde{x})+M(n-d(x,\tilde{x}))}.$ (21)
$P_{\mathrm{Ran-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})-1\mid x,\tilde{x}\right)\geq\frac{1}{M}\cdot\frac{d(x,\tilde{x})}{n}.$ (22)
The following theorem establishes the convergence rates in total variation distance of SOMA, Ran-IMwG, and Sys-IMwG for $n=2.$
@@
For the case $n=2$, under Assumption 2, the convergence rate of the SOMA algorithm is upper bounded by
$r_{\mathrm{SOMA}}\leq\frac{M^{2}+3M-2+\sqrt{M^{4}+2M^{3}+9M^{2}-8M}}{2(1+M)^{2}}.$ (23)
@@
$r_{\mathrm{RAN}}\leq\frac{3M-2+\sqrt{M^{2}+4M-4}}{4M},\quad r_{\mathrm{SYS}}\leq\frac{\sqrt{(M+1)(M-1)}}{M},$ (24)
@@
The proof constructs a transition matrix of the between-chain distance $d(x^{(t)},\tilde{x}^{(t)})$. The hitting time to absorbing state $d=0$ is the coupling time of $P_{\mathrm{SOMA-coupled}}$. See Section B.3 for details.
@@
The analysis we provide in Theorem 12 has focused on $n=2$. It is natural to ask whether Theorem 12 can be extended to the case where $n>2$. Deriving explicit convergence rate upper bounds proves to be difficult for $n\geq 2$. The main challenge lies in controlling of probability of expansion, i.e. $d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1$. While we provide a bound in Lemma 15, the result, when used in conjunction with Lemma 11, does not lead to a ‘contractive coupling’ for $n>2$, in the sense of Blanca et al., (2022, Definition 1.9) and Bubley and Dyer, (1997). In fact, a qualitative comparison between multiple-component Gibbs samplers would also be challenging when $n>2$ and we refer readers to Qin and Jones, (2022, Remark 4.4 and Section 6) for more discussions on the technical hurdles.

### v2

@@ 3 A SINGLE-OFFER-MULTIPLE-ATTEMPTS SAMPLER
SOMA sampler for (3).
@@
   Propose $y\sim q(\cdot)=f(\cdot\mid\theta)$,
   Calculate weights $w_{i}=\eta(s_{\mathrm{dp}}\mid s([x_{-i},y]))$, $w_{0}=\eta(s_{\mathrm{dp}}\mid s(x))$ and $W=\sum_{i=1}^{n}w_{i}$,
   Select random index $I$ from $\{1,2,\ldots,n\}$ with probability $\mathbb{P}(I=i)\propto w_{i}$,
Stage 2: With probability $\alpha=\min\left\{1,\frac{W}{W+w_{0}-w_{I}}\right\}$, replace $x_{I}$ by $y$.
While IMwG exploits the symmetry of permutation invariant distributions by using the same independent proposal for each component, SOMA has pushed this idea further. SOMA recognizes that an unfavorable proposal for $x_{j}$ might be a favorable proposal for some other $x_{i}$, and it seeks to find such an index $i$ by comparing all the weights given in (7).
The advantage of SOMA lies in its efficient use of the ‘single offer’ from the proposal distribution $q$. Stage 1 of SOMA chooses a favorable component $i$ among the multiple choices to swap out $x_{i}$ by $y$. Then Stage 2 compares the current state $x$ with the proposed state $x^{\prime}=[x_{-i},y]$. The acceptance probability (8) accounts for the fact that $x^{\prime}$ is selected from multiple attempts and is not a direct comparison like (5). We summarize SOMA in Algorithm 1 and show that it is reversible with respect to the target $\pi$ in Theorem 5. The proof is deferred to Section A.1.
@@
Assume $\pi$ is permutation invariant. Stage 1 of Algorithm 1 generates a proposal kernel, which we denote by $K^{\mathrm{SOMA}}(x^{\prime}\mid x)$.
  • We have $K^{\mathrm{SOMA}}(x^{\prime}\mid x)=0$ for all $d(x,x^{\prime})>1$.
  • When $d(x,x^{\prime})=1$, suppose they only differ on the $i$-th entry, i.e., $x_{i}^{\prime}=y$ and $x_{j}^{\prime}=x_{j}$ for all $j\neq i$, then $K^{\mathrm{SOMA}}(x^{\prime}\mid x)=q(y){w_{i}(y,x)}/{W(y,x)}.$
  • Associating proposal kernel $K^{\mathrm{SOMA}}$ with acceptance probability (8) leads to a Markov transition kernel $P_{\mathrm{SOMA}}$ reversible with respect to $\pi$.
@@
Theorem 5 shows that SOMA transforms the independent proposal $q$ into an informed proposal kernel $K^{\mathrm{SOMA}}$ that is dependent on the current state. This distinguishes SOMA from IMH.
Corollary 6 (Sufficient conditions for reversibility).
The following conditions are sufficient for reversibility.
  • $\pi$ is a permutation invariant distribution, and
  • The weights are ‘proper’ in the sense
    $w_{0}(y,x)$ $=c\cdot\frac{\pi(x)}{\prod_{j=1}^{n}q(x_{j})},\ \text{and }$
    $w_{i}(y,x)$ $=c\cdot\frac{\pi([x_{-i},y])}{q(y)\prod_{j=1,j\neq i}^{n}q(x_{j})},\ \forall i=1,\ldots,n$
    for some fixed constant $c>0.$
This corollary shows that SOMA can be applied for any permutation invariant target. Applying SOMA for data imputation from privatized summary statistics is a special and important application.
@@
We should also distinguish the SOMA sampler from the Ran-IMwG and Sys-IMwG. To begin with, SOMA is not a Gibbs sampler - it is a Metropolis-Hastings type algorithm, consisting of a proposal step and an accept/reject step. Instead, Stage 1 of SOMA constructs a Metropolis kernel $K^{\mathrm{SOMA}}$ by converting the IMH proposal into a dependent one. Second, although SOMA and Ran-IMwG randomly pick an index $i$ to perform a swap move, Ran-IMwG uses fixed weights independent of the current state, while SOMA uses weights (7) that adapt to the current state. We can see SOMA as a multiproposal MCMC (Pozza and Zanella,, 2025; Lin et al.,, 2025) method, and Ran-IMwG is the ‘single proposal’ counterpart that SOMA should compare with. More discussions are in Section A.3.
@@
Multiple-try Metropolis (MTM) (Liu et al.,, 2000) also entails multiple evaluations in one Metropolis iteration. MTM would require multiple proposed points, independent or correlated (Craiu and Lemieux,, 2007). The Multiple-try Metropolis Independent sampler (MTM-IS) (Yang and Liu,, 2023) bears resemblance to SOMA in using the IMH kernel. MTM-IS generates multiple samples $\{y_{1},\ldots,y_{n}\}$ from an independent proposal $q$, selects one index from them, and generates ‘balancing trials’ to satisfy reversibility. On the other hand, SOMA only requires one sample $y$ from $q$. The multiple attempts for SOMA are created by comparing $y$ with each component in the current state $x$. SOMA would require much less computation and memory cost than MTM-IS.
Remark 4.
Per-iteration cost of SOMA is $\mathcal{O}(n).$ It is natural to modify SOMA to simultaneously consider a subset of $m<n$ components, and thus reducing the cost. Let $M$ be the index set. In Stage 1, one can compute only the weights $w_{0}$ and $w_{i}$ for $i\in M$. In Stage 2, the acceptance probability would be $\min\left(1,\frac{\sum_{i\in M}w_{i}}{w_{0}+\sum_{j\not=I,j\in M}w_{j}}\right).$ Our theoretical analysis considers the standard case, where all $n$ components are considered. The reversibility still holds for this $m$-component variant. See Appendix A.2 for details.
@@ 4.1 Acceptance Rate
For accept-reject-based Markov chains, a small acceptance probability is a clear signal for slow convergence (Brown and Jones,, 2022). For component-wise Markov chains, slow convergence of the Metropolis kernels makes the overall Metropolis within Gibbs sampler mix slowly (Qin et al.,, 2025).
First, we analyze the coordinate-wise acceptance probability of the three algorithms. This is the probability of accepting a proposed move conditioning on the proposed point $y$ as well as the chosen coordinate $i$ for a swap move.
@@
See complete proof in Section A.5.
Theorem 8 indicates that, if all three methods choose to update the same component $i$, SOMA always has an acceptance probability at least as high as the other IMwG methods.
@@
Next, we analyze the overall acceptance probabilities. Let $A(x)$ be the probability of accepting a move from $x$. For SOMA and Ran-IMwG, the coordinate $i$ is chosen randomly, and we have
@@
$\|P^{t}_{\mathrm{RAN}}(\cdot\mid x)-\pi(\cdot)\|_{\mathrm{TV}}$ $\geq\left(1-A^{\mathrm{RAN}}(x)\right)^{t}\geq\left(1-A^{\mathrm{SOMA}}(x)\right)^{t}.$
@@
We now give lower bounds for the overall acceptance probability from the current state $x$ under $\epsilon$-DP.
@@
Let $A^{\mathrm{SOMA}}(x)$, $A^{\mathrm{RAN}}(x)$ and $A_{i}^{\mathrm{SYS}}(x)$ be defined as in (13), (14) and (15) respectively. These are the overall acceptance probabilities from a current state $x$. Under $\epsilon$-DP, we have
$A^{\mathrm{SOMA}}(x)\geq\frac{n}{n+e^{\epsilon}-1},\ A^{\mathrm{RAN}}(x)\geq\frac{1}{e^{\epsilon}},\ \text{and}\ A_{i}^{\mathrm{SYS}}(x)\geq\frac{1}{e^{\epsilon}}\ \forall i.$ (17)
The lower bounds on acceptance probability place the three algorithms into a ‘no clear signal of slow convergence’ scenario. We also highlight that when $n$ is large, the acceptance probability of SOMA approaches 1. As a result, SOMA becomes almost rejection-free in high dimensions.
In fact, our experiments in 5.2 for Bayesian linear regression already reports SOMA acceptance rate between 92% and 99.5% for sample size $n=10$, depending on the privacy loss budget $\epsilon.$
@@ 4.2 Upper Bound of the Convergence Rate
In this section, we use coupling-based techniques to analyze the convergence rate of the algorithms. By constructing a faithful coupling (see Appendix A.6 for technical details and the coupling kernel), we establish the following upper bounds on the convergence rates for the $n=2$ case:
@@
For the case $n=2$, under $\epsilon$-DP, the convergence rate of the SOMA algorithm is upper bounded by
$r_{\mathrm{SOMA}}\leq\frac{e^{2\epsilon}+3e^{\epsilon}-2+\sqrt{e^{4\epsilon}+2e^{3\epsilon}+9e^{2\epsilon}-8e^{\epsilon}}}{2(1+e^{\epsilon})^{2}}.$ (18)
@@
$r_{\mathrm{RAN}}$ $\leq\frac{3e^{\epsilon}-2+\sqrt{e^{2\epsilon}+4e^{\epsilon}-4}}{4e^{\epsilon}},$
$r_{\mathrm{SYS}}$ $\leq\frac{\sqrt{(e^{\epsilon}+1)(e^{\epsilon}-1)}}{e^{\epsilon}},$ (19)
@@
The proof constructs a transition matrix of the between-chain distance $d(x^{(t)},\tilde{x}^{(t)})$. The hitting time to the absorbing state $d=0$ is the coupling time, which yields an upper bound on the convergence rate through the coupling inequality. See Appendix A.6 for details.
@@
These bounds connect the privacy budget $\epsilon$ to sampler efficiency. Interestingly, stronger privacy (smaller $\epsilon$ and larger noise) leads to faster convergence. This occurs as the data imputation posterior becomes more diffusive and hence easier to approximate by IMH using $q(\cdot)=f_{\theta}(\cdot)$. This is also confirmed by our experiments; see Figure 2.
Remark 6.
The analysis we provide in Theorem 13 has focused on $n=2$. It is natural to ask whether Theorem 13 can be extended to the case where $n>2$. Deriving explicit convergence rate upper bounds proves to be difficult for $n\geq 2$. The main challenge lies in controlling the probability of expansion, i.e. $d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1$. While we provide a bound in Lemma 19, the result, when used in conjunction with Lemma 18, does not lead to a ‘contractive coupling’ for $n>2$, in the sense of Blanca et al., (2022, Definition 1.9) and Bubley and Dyer, (1997). In fact, a qualitative comparison between Ran- and Sys-Scan Gibbs would also be challenging when $n>2$, and we refer readers to Qin and Jones, (2022, Remark 4.4 and Section 6) for more discussions on the technical hurdles.